# Existence and qualitative behavior of solutions of abstract differential-algebraic equations

### v1

@@ 4.2 Second approach
Equation (2.20) is equivalent to $\frac{d}{dt}[P_{1}x]+\widetilde{A}^{-1}BP_{1}x=\widetilde{A}^{-1}Q_{1}f(t,x)$, where $\widetilde{A}^{-1}\in\mathrm{L}(Y,D)$ is the inverse of (4.5), or to $\frac{d}{dt}[P_{1}x]+G^{-1}BP_{1}x=G^{-1}Q_{1}f(t,x)$, where $G^{-1}\in\mathrm{L}(Y,D)$ is the inverse of $G=Q_{1}A+Q_{2}B$ [33, Subsection 3.3], since $\widetilde{A}^{-1}AP_{1}=P_{1}$ and $G^{-1}AP_{1}=P_{1}$.
@@
This representation of $x_{2}$ correspond to the direct decomposition $D_{2}=D_{20}\mathop{\dot{+}}D_{2\Sigma}$ (cf. (2.11)).
@@
where $\widehat{\eta}_{20}(t,w_{1}):=\eta_{20}(t,\widetilde{A}^{-1}w_{1})$. Since $\widetilde{A}^{-1}\in\mathrm{L}(Y,D)$, $Q_{1}\in\mathrm{L}(Y)$, $B\widetilde{A}^{-1}\in\mathrm{L}(Y)$, then $\breve{\Pi}\in C({\mathbb{R}}_{+}\times\widetilde{M}_{1},Y_{1})$ and, due to the properties of $Q_{1}f$ and $\eta_{20}$, it satisfies locally a Lipschitz condition with respect to $w_{1}$ on ${\mathbb{R}}_{+}\times\widetilde{M}_{1}$. Using [29, p. 9, Theorem 1], we infer that for each initial point $(t_{0},x_{0,1})\in{\mathbb{R}}_{+}\times M_{1}$ the IVP for equation (4.58) with the initial condition
@@
has a unique solution $w_{1}(t)$ on some interval $J\subset{\mathbb{R}}_{+}$ that contains $t_{0}$. Further, by [29, p. 16-17, Corollary] we obtain that the solution $w_{1}(t)$ can be extended over a maximal interval of existence $t_{0},t_{max})$  ($t_{max}\leq\infty$) of $w_{1}(t)$ in $\widetilde{M}_{1}$, and the extended solution is unique. Notice that the function $x(t)=\widetilde{A}^{-1}w_{1}(t)+\widehat{\eta}_{20}(t,w_{1}(t))+\eta_{2\Sigma}(t)$ is continuous and $Ax(t)=w_{1}(t)+A\eta_{2\Sigma}(t)$ has the continuous derivative $\dfrac{dAx}{dt}(t)=\breve{\Pi}(t,w_{1}(t))+AP_{2\Sigma}\dfrac{d\eta_{2\Sigma}}{dt}(t)$ on $t_{0},t_{max})$ (for $t_{0}=0$, $Ax(t)$ has the derivative on the right at $t_{0}$). Thus, the IVP (1.1), (4.38), (4.45) has the unique solution $x(t)=\widetilde{A}^{-1}w_{1}(t)+\widehat{\eta}_{20}(t,w_{1}(t))+\eta_{2\Sigma}(t)$ in $M$ on the maximal interval of existence $t_{0},t_{max})$. ∎
@@
The operator $\mathcal{B}_{2}^{(-1)}\in\mathrm{L}(Y,D)$} is defined by the relations $\mathcal{B}_{2}^{(-1)}\mathcal{B}_{2}=P_{2}$}, $\mathcal{B}_{2}\mathcal{B}_{2}^{(-1)}=Q_{2}$} and $\mathcal{B}_{2}^{(-1)}=P_{2}\mathcal{B}_{2}^{(-1)}$}.
@@ 4.3 Cases of the characteristic pencil of index 1 and 2
$D_{20}=D_{20}^{(1)}\mathop{\dot{+}}D_{20}^{(2)},\quad D_{20}^{(2)}=D_{2\wedge,0},\quad D_{21}=D_{2\Sigma}=D_{2\Sigma}^{(1)},$
$Y_{20}=Y_{20}^{(1)}\mathop{\dot{+}}Y_{20}^{(2)},\quad Y_{20}^{(2)}=Y_{2\Sigma}=Y_{2\Sigma,0},\quad Y_{20}^{(1)}=Y_{2*}^{(1)},\quad Y_{21}=Y_{21}^{(2)}=Y_{2*}^{(2)}.$
@@
First, we specify the results of Section 4.1. The ADAE (1.1) with the regular pencil $P(\lambda)$ of index 2 is equivalent to the system (4.3), (4.4), where $x_{2\Sigma}=x_{21}$}, $Q_{2*}=Q_{20}^{(1)}+Q_{21}$} and $Q_{2\Sigma}=Q_{20}^{(2)}$}. In addition, for the theorems and other results of Section 4.1 we have $D_{2\Sigma}=D_{21}$}, $Y_{2*}=Y_{20}^{(1)}\mathop{\dot{+}}Y_{21}$}, $Y_{2\Sigma}=Y_{20}^{(2)}$}, $P_{2\Sigma}=P_{21}$}, and it is convenient to denote $M_{21}:=M_{2\Sigma}$}. Taking into account the mentioned relations, it is easy to formulate the results (theorems, corollaries, etc.) of Section 4.1 in an appropriate way.
Now let us specify the results of Section 4.2. The function $f$ having the structure (4.38) as well as (4.76) (recall that by construction $x_{21}=x_{21}^{(2)}$ and $Q_{21}=Q_{21}^{(2)}$) takes the form

### v2

@@ 4.2 Second approach
As mentioned in Remark 4.1, for convenience we assume that $D_{B}\supseteq D_{A}$.
Equation (2.20) is equivalent to $\frac{d}{dt}[P_{1}x]+\widetilde{A}^{-1}BP_{1}x=\widetilde{A}^{-1}Q_{1}f(t,x)$, or $\frac{d}{dt}[P_{1}x]+A^{(-1)}BP_{1}x=A^{(-1)}Q_{1}f(t,x)$, or $\frac{d}{dt}[P_{1}x]+G^{-1}BP_{1}x=G^{-1}Q_{1}f(t,x)$, where $\widetilde{A}^{-1}\in\mathrm{L}(Y,X)$ is the inverse of (4.5), $A^{(-1)}\in\mathrm{L}(Y,X)$ is the semi-inverse of $A$, and $G^{-1}\in\mathrm{L}(Y,X)$ is the inverse of $G=Q_{1}A+Q_{2}B$ [32, Subsection 3.3], since $\widetilde{A}^{-1}AP_{1}=A^{(-1)}AP_{1}=G^{-1}AP_{1}=P_{1}$.
@@
This representation of $x_{2}$ correspond to the direct decomposition (cf. (2.11))
$D_{2}=D_{20}\mathop{\dot{+}}D_{2\Sigma}.$ (4.39)
@@
where $\widehat{\eta}_{20}(t,w_{1}):=\eta_{20}(t,\widetilde{A}^{-1}w_{1})$. Since $\widetilde{A}^{-1}\in\mathrm{L}(Y,X)$, $Q_{1}\in\mathrm{L}(Y)$, $B\widetilde{A}^{-1}\in\mathrm{L}(Y)$, then $\breve{\Pi}\in C({\mathbb{R}}_{+}\times\widetilde{M}_{1},Y_{1})$ and, due to the properties of $Q_{1}f$ and $\eta_{20}$, it satisfies locally a Lipschitz condition with respect to $w_{1}$ on ${\mathbb{R}}_{+}\times\widetilde{M}_{1}$. Using [28, p. 9, Theorem 1], we infer that for each initial point $(t_{0},x_{0,1})\in{\mathbb{R}}_{+}\times M_{1}$ the IVP for equation (4.60) with the initial condition
@@
has a unique solution $w_{1}(t)$ on some interval $J\subset{\mathbb{R}}_{+}$ that contains $t_{0}$. Further, by [28, p. 16-17, Corollary] we obtain that the solution $w_{1}(t)$ can be extended over a maximal interval of existence $t_{0},t_{max})$  ($t_{max}\leq\infty$) of $w_{1}(t)$ in $\widetilde{M}_{1}$, and the extended solution is unique. Notice that the function $x(t)=\widetilde{A}^{-1}w_{1}(t)+\widehat{\eta}_{20}(t,w_{1}(t))+\eta_{2\Sigma}(t)$ is continuous and $Ax(t)=w_{1}(t)+A\eta_{2\Sigma}(t)$ has the continuous derivative $\dfrac{dAx}{dt}(t)=\breve{\Pi}(t,w_{1}(t))+AP_{2\Sigma}\dfrac{d\eta_{2\Sigma}}{dt}(t)$ on $t_{0},t_{max})$ (for $t_{0}=0$, $Ax(t)$ has the derivative on the right at $t_{0}$). Thus, the IVP (1.1), (4.40), (4.47) has the unique solution $x(t)=\widetilde{A}^{-1}w_{1}(t)+\widehat{\eta}_{20}(t,w_{1}(t))+\eta_{2\Sigma}(t)$ in $M$ on the maximal interval of existence $t_{0},t_{max})$. ∎
@@
The operator $\mathcal{B}_{2}^{(-1)}\in\mathrm{L}(Y,X)$} is defined by the relations $\mathcal{B}_{2}^{(-1)}\mathcal{B}_{2}=P_{2}$}, $\mathcal{B}_{2}\mathcal{B}_{2}^{(-1)}=Q_{2}$} and $\mathcal{B}_{2}^{(-1)}=P_{2}\mathcal{B}_{2}^{(-1)}$}.
@@
Remark 4.10.
As in Section 4.1 (see Remark 4.7), in the above theorems, lemmas and corollaries one can use the semi-inverse operator $A^{(-1)}$ defined by (4.6) instead of the inverse $\widetilde{A}^{-1}$ of (4.5), taking into account that $A^{(-1)}=\widetilde{A}^{-1}(Q_{1}+Q_{2\Sigma})$},  $\widetilde{A}^{-1}\widetilde{M}_{1}=A^{(-1)}\widetilde{M}_{1}$}  ($\widetilde{M}_{1}\subseteq Y_{1}$),  $\widetilde{A}x_{1}=Ax_{1}$,  $\widetilde{A}^{-1}w_{1}=A^{(-1)}w_{1}$  ($w\in Y_{1}$),  $\widetilde{A}^{-1}\widetilde{\Omega}_{1}=A^{(-1)}\widetilde{\Omega}_{1}$  ($\widetilde{\Omega}_{1}\subseteq Y_{1}$).
@@ 4.3 Cases of the characteristic pencil of index 1 and 2
$\begin{split}&D_{20}=D_{20}^{(1)}\mathop{\dot{+}}D_{20}^{(2)},\quad D_{20}^{(2)}=D_{2\wedge,0},\quad D_{21}=D_{2\Sigma}=D_{2\Sigma}^{(1)},\\ &Y_{20}=Y_{20}^{(1)}\mathop{\dot{+}}Y_{20}^{(2)},\quad Y_{20}^{(2)}=Y_{2\Sigma}=Y_{2\Sigma,0},\quad Y_{20}^{(1)}=Y_{2*}^{(1)},\quad Y_{21}=Y_{21}^{(2)}=Y_{2*}^{(2)}.\end{split}$ (4.81)
@@
First, we specify the results of Section 4.1.
The ADAE (1.1) with the regular pencil $P(\lambda)$ of index 2 is equivalent to the system (4.3), (4.4) (or (4.1), (4.2)), where $x_{2\Sigma}=x_{21}$}, $Q_{2*}=Q_{20}^{(1)}+Q_{21}$} and $Q_{2\Sigma}=Q_{20}^{(2)}$}, i.e.,
$\frac{d}{dt}[x_{1}+x_{21}]=\Pi(t,x),\quad\Pi(t,x):=\widetilde{A}^{-1}(Q_{1}+Q_{20}^{(2)})\big[f(t,x)-Bx\big],$ (4.82)
$F_{2*}(t,x_{1},x_{21},x_{20})=0,\quad F_{2*}(t,x_{1},x_{21},x_{20}):=(Q_{20}^{(1)}+Q_{21})[f(t,x)-B(x_{21}+x_{20})].$ (4.83)
In addition, for the theorems and other results of Section 4.1 we have $D_{2\Sigma}=D_{21}$}, $Y_{2*}=Y_{20}^{(1)}\mathop{\dot{+}}Y_{21}$}, $Y_{2\Sigma}=Y_{20}^{(2)}$}, $P_{2\Sigma}=P_{21}$}, and it is convenient to denote $M_{21}:=M_{2\Sigma}$}. Recall that in Section 4.1 we consider the case when the component $x_{20}$ can be defined as an implicit function from (4.4), thus the function $F_{2*}$ should depend on $x_{20}$. Equation (4.4) takes the form (4.83) for the ADAE with the index-2 regular pencil. Taking into account the mentioned relations, it is easy to formulate the results (theorems, corollaries, etc.) of Section 4.1 in an appropriate way.
Now let us specify the results of Section 4.2.
The function $f$ having the structure (4.40) as well as (4.78) (recall that by construction $x_{21}=x_{21}^{(2)}$ and $Q_{21}=Q_{21}^{(2)}$) takes the form
@@
5 Applications
5.1 Example
Consider the initial boundary value problem (IBVP) from [26, Section 5.2]:
$\begin{split}&\partial_{t}\partial^{2}_{x}u+\left(\partial^{2}_{x}+1\right)u=f(t,u)\\ &2u(t,1)-2u(t,0)-\partial_{x}u(t,1)=0,\\ &\partial_{x}u(t,0)=0,\\ &u(0,x)=0,\end{split}$ (5.1)
where $t\in{\mathbb{R}}_{+}$, $x\in[0,1]$. It can be represented as the following IVP for the ADAE of the type (1.1) in $L_{2}[0,1]$:
$\begin{split}&\frac{d}{dt}[Au]+Bu=f(t,u),\quad A=\partial^{2}_{x},\quad B=\partial^{2}_{x}+1,\\ &u(0)=0,\end{split}$ (5.2)
where $u\in L_{2}[0,1]$ and $A$, $B$ are defined on $D_{A}=D_{B}=D=\{{y\in L_{2}[0,1]}\mid y^{\prime}(x)\text{ is absolutely continuous on $[0,1]$},\;y^{\prime}(0)=0,\;2y(1)-2y(0)-y^{\prime}(1)=0\}$.
The resolvent $(A+\mu B)^{-1}$ has a pole of order 2 at the point $\mu=0$. Therefore, $P(\lambda)$ is a regular pencil of index 2. Since $\ker A=\mathop{\mathrm{span}}\{\varphi_{1}^{1}\}$, where $\varphi_{1}^{1}=1$, and the index $\nu=2$, then the canonical system of root vectors of the pencil $A+\mu B$ at $\mu=0$ consists of the vectors $\varphi_{1}^{1}$, $\varphi_{1}^{2}$, where $\varphi_{1}^{2}$ can be found from the relation $A\varphi_{1}^{2}=-B\varphi_{1}^{1}$ (see (2.6)). The projectors take form
$P_{20}u=(u,p^{1}_{1})\varphi_{1}^{1},\quad P_{21}u=(u,p_{1}^{2})\varphi_{1}^{2},\quad P_{2}=P_{20}+P_{21},\quad P_{1}=I_{X}-P_{2},$
$Q_{20}v=(v,q_{1}^{1})B\varphi_{1}^{1},\quad Q_{21}v=(v,q_{1}^{2})B\varphi_{1}^{2},\quad Q_{2}=Q_{20}+Q_{21},\quad Q_{1}=I_{Y}-Q_{2},$
where
$\varphi_{1}^{1}=1,\quad\varphi_{1}^{2}=\sqrt{5}x^{2},$
$p_{1}^{1}=B^{*}q_{1}^{1}=4x^{3}-6x^{2}-\frac{8}{5}\left(x-\frac{1}{2}\right)+2,$
$p_{1}^{2}=B^{*}q_{1}^{2}=\frac{12}{\sqrt{5}}\left(x-\frac{1}{2}\right)=q_{1}^{2},$
$B\varphi_{1}^{1}=1,\quad B\varphi_{1}^{2}=\sqrt{5}(x^{2}+2),$
$q_{1}^{1}=4x^{3}-6x^{2}-\frac{128}{5}\left(x-\frac{1}{2}\right)+2,$
$q_{1}^{2}=\frac{12}{\sqrt{5}}\left(x-\frac{1}{2}\right).$
and the subspaces from the direct decompositions (2.10), (2.11), (4.39), (4.81) take the form $D_{20}=\mathop{\mathrm{span}}\{\varphi_{1}^{1}\}$, $D_{21}=D_{2\Sigma}=\mathop{\mathrm{span}}\{\varphi_{1}^{2}\}$, $Y_{20}=Y_{2\Sigma}=\mathop{\mathrm{span}}\{B\varphi_{1}^{1}\}$, and $Y_{21}=Y_{2*}=\mathop{\mathrm{span}}\{B\varphi_{1}^{2}\}$.
Using the relations (4.6), we obtain that the semi-inverse operator of $A$ has the form $A^{(-1)}y(v)=\int\limits_{0}^{v}\int\limits_{0}^{s}y(\zeta)\,d\zeta\,ds-(v^{4}+12v^{2})\int\limits_{0}^{1}\left(s-\frac{1}{2}\right)y(s)\,ds$. Thus, the inverse of (4.5) has the form $\widetilde{A}^{-1}=A^{(-1)}+\langle\cdot,q_{1}^{2}\rangle\varphi_{1}^{1}$.
Further, verifying the conditions of the theorems proved above, we find the requirements necessary for the global solvability and Lagrange stability of the ADAE (5.1), or the conditions under which the solutions blow up in finite time.
Acknowledgments
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, ERC Advanced Grant NEUROMORPH, no. 101018153.